find two numbers whose sum is 26,if the sum or their squares is 340.

Answer by lwsshak3(11628)

find two numbers whose sum is 26,if the sum or their squares is 340.
***
let x=one of the numbers
26-x=other number
x^2+(26-x)^2=340
x^2+676-52x+x^2=340
2x^2-52x+336=0
x^2-26x+168=0
(x-12)(x-14)=0
x=12
or
x=14
..
one of the numbers=12
other number=14